OpenStudy (anonymous):

Will fan and medal!!!! (a + 3)^2/a - 3 divided by 5/(a^2 - 9)

OpenStudy (campbell_st):

when you divide by a fraction, the rule is flip the 2nd fraction and multiply so the problem becomes \[\frac{(a + 3)^2}{(a -3)} \times \frac{(a^2 - 9)}{5}\] factor the quadratic, which is the numerator of the 2nd factor and you'll see a common factor that can be cancelled... and then its an easy simplification. hope it helps